1. Q2(b)(ii) A monopolist has a production function ?=?(?,?)=2∗(??)0.5 with ?=?=1 (input markets are competitive, as normal). For parts (a)-(c), she faces a single-price demand curve ?=5−?.

                b.(ii) What is the monopolist’s profit?
Answer #1

Hi! Welcome to Chegg!

2.

b.

We have production function:

Cost = wl + vk = l + k

Cost minimizing condition is:

MRTS = Input price ratio

gives: k = l

Using this in production function, we get:

Cost function can be calculated as:

Marginal cost is given as:

We have Demand:

p = 5 - q

Profit maximizing condition: MR = MC

5 - 2q = 1

gives: q = 2

Profit maximizing quantity = 2.

Profit is given as:

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!